# BAGELS: Benchmarking the Automated Generation and Extraction of Limitations from Scholarly Text

### v1

@@ 9 Conclusion
We present a new approach for automatically extracting, generating and evaluating limitations in scientific articles. Our method explores chunk analysis, accommodating top sections of the entire paper and integrating OpenReview feedback to capture perspectives beyond those of the original authors. To evaluate the effectiveness of our system, we introduce a granular text evaluation framework that breaks down limitations into more minor points and employs LLMs as a Judge for assessing topic-level alignment. Combining standard NLP metrics with a keyword-based analysis provides a more nuanced and reliable assessment. Our findings are further validated by human review, which reveals a strong alignment between human judgments and the LLM-driven extraction process, including when the LLM is used as a judge.
@@
In this work, we focused on venues in natural language processing (ACL papers from 2023-2024) and machine learning (NeurIPS papers 2021-2022), and Biology domain papers from PeerJ, which ensures high relevance and quality but insufficient for broader generalizability. While this scope allows us to benchmark the performance of LLMs in extracting limitations from well-structured scientific texts, we acknowledge that the findings may not generalize to papers from other fields, such as social sciences, physics, chemistry, or mathematics where writing conventions and limitation styles may differ. Due to high API costs, we did not experiment with GPT-4 or GPT-4o; instead, we opted for GPT-4o Mini as a cost-effective alternative. Similarly, to avoid the high expense of maintaining an extensive vector database for RAG with closed LLMs, we limited our retrieval corpus to 100 randomly selected papers. We did not evaluate Open LLMs with RAG, such as Llama and Mistral, to assess their performance in retrieval-augmented settings. While we incorporated OpenReview comments for NeurIPS papers, we could not find them for ACL papers. Furthermore, we relied only on GPT-4o Mini as the evaluation judge and did not experiment with other LLMs for assessment. To evaluate the effectiveness of LLMs as both text extractors and judges, we conducted a human annotation study with 200 samples and only three annotators.
We used GPT 4o mini for text extraction, text generation, and LLM as a judge. We experimented with other LLMs such as Llama but GPT 4o mini performs better. To rely on LLM for limitation extraction and evaluation ther are some potential biases.
  • Self-Validation Bias: We used the same LLM for text generation and evaluation in experiments. Where a model can favor its style or logic.
  • Positional bias: For LLM as a judge, we have multiple texts to check whether both are similar. Here, LLM can have positional bias, which is more centered on the first input text.
  • Confirmation bias: We experimented with the same LLM (GPT 4o mini) for text extraction and generation. Even though both are used in an isolated environment with zero shot (no fine-tuning), the same LLM can produce confirmation bias on some common/generic limitations such as “Generalizability issue" or “small sample size" which reduces novelty.
  • Hallucination and Overconfidence: When extracting limitations, LLM might sometimes invent limitations that aren’t grounded in the source text.
To mitigate self-validation bias, we compared our LLM’s judgments against human evaluation, tested alternative models for generation and judgment (Llama and Mistral), and incorporated RAG. We changed the order and considered the consistent one to overcome the positional bias problem. Moreover, we integrated RAG with LLM for text generation, where LLM can get diverse evidence or force diversity sampling on generations to minimize the confirmation bias problem. To overcome hallucination and overconfidence, we evaluate the extraction process by LLM with three annotators to determine whether LLM extracts limitations without inventing anything new. However, these approaches are not sufficient.
As part of our future work, we plan to expand the dataset to include papers from a broader range of domains, such as arXiv submissions from diverse subject areas (e.g., bioinformatics, cognitive science) and other peer-reviewed venues beyond ACL and NeurIPS. We also aim to incorporate more review texts from platforms, where available, across different conferences. This expansion will allow us to test the robustness of our limitation extraction models across domains and improve generalizability. We will explore broader contexts from open LLMs incorporating RAG, and integrating more extensive multi-agent LLM frameworks. Moreover, we will gather annotations from a larger and more diverse pool of experts, enabling a deeper analysis of generated limitations. In doing so, we aim to make our framework more robust and broadly applicable while maintaining its core strength of providing granular, interpretable assessments of LLM-generated limitations.
@@
This research adheres to the ethical standards in scientific research. All data used in this study, including research papers and OpenReview feedback, were obtained from publicly available sources and used in compliance with their respective usage policies. No private or sensitive information was accessed or utilized. We acknowledge the potential biases inherent in large language models (LLMs) and have taken steps to mitigate their impact. For instance, we conducted human evaluations to validate the outputs of LLMs, ensuring alignment with human judgment and reducing the risk of generating misleading or inaccurate content. Additionally, our evaluation framework emphasizes transparency and granularity, enabling a more nuanced assessment of generated limitations. For the user study, we engaged three computer science graduate students; they are no longer involved in this research. Their participation was voluntary, and they had no conflicts of interest regarding the study. We are committed to promoting the responsible use of AI in scientific research. Our work aims to enhance scientific studies’ transparency, reliability, and reproducibility by generating and evaluating limitations systematically and ethically. We encourage further research to address potential biases in LLMs and improve their application in scientific contexts. Our study didn’t filter out any dataset based on race, ethnicity, color, location, gender, or other discriminatory attributes. Therefore, we are dedicated to complying with the ethical code and policy attributed to ACL. Large language models inherently carry biases and risks derived from their training corpora—confirmation bias toward “safe” limitations (e.g., small sample size), fluency bias favoring well-written over factually grounded outputs, verbosity bias rewarding longer responses, and self-validation bias when the same model is used for extraction, generation, and evaluation. To mitigate these, we ground all generations in source content and peer reviews via Retrieval-Augmented Generation (RAG), diversify our ground truth by incorporating human-authored OpenReview critiques, and conduct parallel human evaluations to detect overconfidence and evaluation bias. RAG-based outputs are demonstrably more concise, reducing verbosity bias, and our multi-model judge setup (including Llama and Mistral alongside GPT-4o mini) helps break circularity. We acknowledge that further work is needed to quantify these biases rigorously and plan to investigate cross-domain robustness and fairness in future studies.
@@ A.2 Performance Measurement
We used three distinct prompts to check the topic level similarity between ground truth limitations and LLM-generated limitations (Figure 3, Appendix). To overcome the position bias, we choose the consistent one.
Metric Limitation Limitaiton + OPR
Rouge-L 14.59 14.52 (-0.07)
BERTScore 84.59 84.54 (-0.05)
JStext 7.70 7.62 (-0.08)
CStext 30.95 32.02 (+1.07)
CGT 38.28 32.51 (-5.77)
CLLM 22.09 44.46 (+22.37)
F1 score 0.24 0.34 (+0.1)
Table 7: Ablation study with Mistral 7b in NeurIPS data, considering Limitation and Limitation + OpenReview (OPR) as ground truth. Top 3 sections are input here.
We also compared BERTScore between the Traditional approach vs the Pointwise approach, where the Pointwise approach preserves nuanced details and showed better performance (Table 8, Appendix).
Approach Traditional way Pointwise approach
T5 82.60 87.81 (+5.21)
BART 83.09 87.67 (+ 4.58)
Pegasus 82.76 87.66 (+4.9)
GPT 3.5 83.79 87.08 (+3.29)
GPT 4o mini 82.66 86.02 (+3.36)
Table 8: Comparison of Traditional Approach and Pointwise Approach. BERTScore of each model in ACL data (Limitation is ground truth)
@@
Model R-1 R-2 R-L BertScore BertScoreheading JStext CStext BLEU CSkeyword JSkeyword
T5 27.13 5.11 19.92 87.81 - 10.82 31.79 0.4 18.69 11.26
BART 26.05 5.49 19.43 87.67 - 10.68 31.91 0.7 19.89 12.19
Pegasus 27.0 5.46 20.15 87.66 - 10.71 33.39 0.6 17.65 10.58
GPT 3.5 32.74 10.92 24.24 87.08 85.50 14.65 43.12 3.54 26.54 17.46
Llama 3 33.98 12.02 25.66 88.30 85.30 14.69 40.4 3.6 30.33 20.28
GPT 4o-mini 22.90 3.99 16.57 86.02 75.02 8.70 32.29 0.33 16.20 9.85
GPT 3.5 + RAG 38.11 16.31 30.21 90.88 87.65 19.47 45.37 6.64 35.74 24.65
GPT 4o mini + RAG 31 9.95 23.17 87.33 - 12.99 39.29 2.69 26.47 17.19
Table 10: Results of models in “performance” metrics – Rouge-x, BLEU, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading (ACL data).
Metric ACL NeurIPS PeerJ
Rouge-L 16.57 17.07 10.62
BERTScore 86.02 86.25 83.26
JStext 8.70 9.12 5.81
CStext 32.29 35.42 26.19
CSkeyword 16.20 15.67 2.89
JSkeyword 9.85 9.31 1.66
BLEU 0.33 0.4 0.01
CGT 57.65 45.69 55.07
CLLM 19.76 76.27 35.87
F1 score 0.24 0.54 0.39
Table 11: Comparison among ACL and NeurIPS PeerJ data in terms of Coverage and Performance. Full text is input, using GPT 4o mini to generate limitations, and Ground truth is Limitation + Peer Review (for NeurIPS and PeerJ), ACL has only author-mentioned Limitations.
Metric Input Ground Truth
Sec Auth OPR Auth + OPR
R-1 3 22.90 26.13 23.69 ( +0.79)
R-1 All 23.74 (
$\uparrow$
26.36 (
$\uparrow$
26.36 (+2.62)
R-2 3 3.99 5.44 4.75
(+0.76)
R-2 All 4.51 (
$\uparrow$
5.41 (
$\downarrow$
5.64
(+1.13)
R-L 3 16.57 18.55 17.07
(+0.5)
R-L All 17.11 (
$\uparrow$
18.65 (
$\uparrow$
18.83
(+1.72)
BS 3 86.02 86.50 86.25
(+0.23)
BS All 86.24 (
$\uparrow$
85.96 (
$\downarrow$
86.67
(+0.43)
BSheading 3 75.02 85.79 69.50
(-5.52)
BSheading All 85.37 (
$\uparrow$
85.75 (
$\downarrow$
85.59
(+0.22)
JStext 3 8.70 10.17 9.12
(+0.42)
JStext All 9.22 (
$\uparrow$
10.23 (
$\uparrow$
10.38
(+1.16)
CStext 3 32.29 38.35 35.42
(+3.13)
CStext All 33.23 (
$\downarrow$
38.74 (
38.87
(+5.64)
BLEU 3 0.33 0.5 0.4
(+0.07)
BLEU All 0.4 (
$\uparrow$
0.5 (
$\uparrow$
0.6
(+0.2)
CSKeyword 3 16.20 15.15 15.67
(-0.53)
CSKeyword All 18.14 (
$\uparrow$
15.42 (
$\uparrow$
16.83
(-2.72)
JSKeyword 3 9.85 8.99 9.31
(-0.54)
JSKeyword All 11.10 (
$\uparrow$
9.21 (
$\uparrow$
10.19
(-0.91)
CGT 3 57.65 54.79 45.69
(-12.41)
CGT All 61.55 (
$\uparrow$
56.55 (
$\uparrow$
49.43
(-12.12)
CLLM 3 19.76 56.79 76.27
(+57.07)
CLLM All 40.04 (
$\uparrow$
56.77 (
$\downarrow$
42.59
(+2.55)
Pr. 3 0.20 0.56 0.65
(+0.46)
Pr. All 0.4 (
$\uparrow$
0.57 (
$\uparrow$
0.43
(+0.03)
Rec. 3 0.31 0.55 0.46
(+0.15)
Rec. All 0.51 (
$\uparrow$
0.56 (
$\uparrow$
0.46
(-0.05)
F1 3 0.24 0.55 0.54
(+0.3)
F1 All 0.45 (
$\uparrow$
0.56 (
$\uparrow$
0.44
(-0.01)
Table 12: GPT 4o mini results in “coverage” (Coverage of Ground Truth Limitation (CGT), LLM Generated Limitation (CLLM), Precision, Recall, and F1-score) and “performance” (Rouge-x, BertScore (over entire text and just heading), Jaccard (JS) and Cosine (CS) similarity over text and heading) metrics when different input sizes and sources are used for limitation extraction. Increasing input length and including external reviews improve generation results (NeurIPS data).
Metric GPT 4o mini GPT 4o mini + RAG
Rouge-1 23.69 24.69 (+1.0)
Rouge-2 4.75 5.24 (+0.49)
Rouge-L 17.07 18.10 (+1.03)
BERTScore 86.25 86.45 (+0.2)
BSheading 69.50 74.74 (+5.24)
JStext 9.12 9.72 (+0.6)
CStext 35.42 34.71 (-0.71)
CSkeyword 15.67 15.86 (+0.19)
JSkeyword 9.31 9.60 (+0.29)
BLEU 0.4 0.7 (+0.3)
CGT 45.69 61.76 (+16.07)
CLLM 76.27 59.18 (-17.09)
Precision 0.65 0.60 (-0.05)
Recall 0.46 0.59 (+0.13)
F1 score 0.54 0.60 (+0.06)
Table 13: Coverage and Performance measure using GPT 4o mini + RAG in NeurIPS data. Ground truth is Limitation + OpenReview, and input is the top 3 sections alongside the abstract.
Metric Input (3 sections) Input (All Sections)
Rouge-1 21.32 20.5 (-0.82)
Rouge-2 3.91 4.0 (+0.09)
Rouge-L 14.59 13.95 (-0.64)
BERTScore 84.59 84.23 (-0.36)
JStext 7.70 7.49 (-0.21)
CStext 30.95 31.82 (+0.87)
BLEU 0.3 0.3 (0)
CSkeyword 16.43 17.29 (+0.86)
JSkeyword 10.00 10.59 (+0.59)
CGT 38.28 37.99 (-0.29)
CLLM 22.09 30.44 (+8.35)
F1 score 0.24 0.32 (+0.08)
Table 14: Ablation study with Mistral 7b in Neurips data. Considering top 3 sections and all sections as input. Author mentioned limitation is ground truth here.
Metric Input (3 sections) Input (All Sections)
Rouge-1 33.98 29.57 (-4.41)
Rouge-2 12.02 9.43 (-2.59)
Rouge-L 25.66 22.78 (-2.88)
BERTScore 88.30 87.29 (-1.01)
JStext 14.69 11.16 (-3.53)
CStext 40.40 34.80 (-5.6)
BLEU 3.61 1.22 (-2.39)
CSkeyword 30.33 30.73 (+0.4)
JSkeyword 20.28 20.31 (+0.03)
CGT 61.38 18.01 (-43.37)
CLLM 39.04 5.01 (-34.03)
F1 score 0.44 0.05 (-0.39)
Table 15: Ablation study with Llama 3 8b in ACL data. Considering top 3 sections and All sections as input. Author mentioned limitation is ground truth here.
Metric Limitation Limitaiton + OpenReview
Rouge-1 21.32 21.02 (-0.3)
Rouge-2 3.91 3.74 (-0.17)
Rouge-L 14.59 14.52 (-0.07)
BERTScore 84.59 84.54 (-0.05)
JStext 7.70 7.62 (-0.08)
CStext 30.95 32.02 (+1.07)
BLEU 0.3 0.2 (-0.1)
CSkeyword 16.43 14.69 (-1.74)
JSkeyword 10.00 8.76 (-1.24)
CGT 38.28 32.51 (-5.77)
CLLM 22.09 44.46 (+22.37)
F1 score 0.24 0.34 (+0.1)
Table 16: Ablation study with Mistral 7b in NeurIPS data, considering Limitation and Limitation + OpenReview as ground truth. Top 3 sections are input here.
Approach Limitation OpenReview Limitaiton + OpenReview
Mistral (Traditional) 70.67 70.19 70.38
Mistral (Pointwise) 84.23 (+13.56) 84.59 (+14.4) 84.54 (+14.16)
Llama (Traditional) 67.55 67.17 67.50
Llama (Pointwise) 87.29 (+19.74) - -
Table 17: BERTScore Comparison Between PointWise (Proposed) and Traditional Approaches Across three different Ground Truths (Limitaiton, OpenReview, Limitation + OpenReview) in NeurIPS Data
Agreement User 1 User 2 User 3
User 1 1 0.75 0.38
User 2 - 1 0.55
User 3 - - 1
Table 18: Inter annotator agreement amont human annotator for GPT 4o mini as Extractor.
@@
Model Name Top 3 Sections All Sections
CGT CLLM F1 Score CGT CLLM F1 Score
Llama 3 61.38 39.04 0.44 18.01 (-43.37) 5.01 (-34.03) 0.05 (-0.39)
GPT-4o Mini 57.65 19.76 0.24 61.55 (+3.9) 40.28 (+20.52) 0.46 (+0.22)
Mistral 38.28 22.09 0.24 37.99 (-0.29) 30.44 (+8.35) 0.32 (+0.08)
Table 20: Performance (Coverage) of various models considering the top three sections versus the entire paper.
Note: CGT, CLLM denotes Coverage of Ground Truth Limitation, Coverage of LLM Generated Limitation. The ’+’ sign in parentheses indicates the performance gain achieved by using all sections instead of only the top three sections.
@@
Figure 3: LLM as a Judge for each limitation.
Note: We use three distinct prompts to verify consistency.
Ground Truth Limitation LLM Generated Limitation BS BS(h) Match
Error-Prone Models:
- Evaluation at the grade school mathematics level indicates that LLMs are still error-prone, but this does not provide insight into their effectiveness in tutoring other subjects.
Keywords: tutoring, LLMs, mathematics, models, subjects
Reliance on Large Language Models (LLMs): The research paper’s dependence on LLMs as Intelligent Tutoring Systems (ITS) for elementary school mathematics may limit the generalizability of the findings.
Keywords: tutoring, mathematics, elementary, models, language
87.88 81.08 Yes
Errors in Code Generation: The presence of significant errors in code generation indicates that LLMs still face challenges in accurately performing tutoring tasks.
Keywords: tutoring, LLMs, errors, code, challenges
90.05 81.38 Yes
Table 21: Ground Truth Limitation matched with multiple LLM Generated Text in PointWise Evaluation.
Note: BS, BS(h), Match denotes BERTScore, BERTSCore (heading), and LLM as a judge decision.
Prompt 1 = ''' Here is the text containing extracted limitations. Please identify and list each limitation, ensuring that each one addresses a distinct topic or point. '''
Figure 4: Prompt to extract limitations from ground truth text.
Ground Truth Limitation LLM Generated Limitation BS BS(h) Match
Quality of Representations vs. Range of Words: The primary limitation is not the quality of the word representations produced, but rather the restricted range of words that the methodologies can be applied to.
Keywords: representations, words, limitation, methodologies, range
Dictionary Limitations: Dictionaries may not include definitions for all word forms in these languages, which restricts the applicability of the proposed approach.
Keywords: dictionaries, dictionary, languages, forms, definitions
89.12 83.08 Yes
Dependence on Bilingual Dictionaries: The approach relies on bilingual dictionaries, which may not cover all necessary vocabulary or nuances of the languages.
Keywords: bilingual, dictionaries, languages, vocabulary, relies
91.71 82.78 Yes
Table 22: LLM Generated Limitation matched with multiple Ground Truth Limitation.
Note: BS, BS(h), Match denotes BERTScore, BERTSCore (heading), and LLM as a judge decision.
@@
**Model Complexity Concerns**: - There is a question regarding whether the performance gains are due to increased model complexity rather than the proposed recursive mixing approach. **Potential for Increased Complexity**: Although the method is described as simple, the recursive nature of the approach may introduce complexity in implementation and understanding, particularly for practitioners who may not be familiar with the underlying concepts Yes Yes Yes
@@
**Computational Complexity**: The Dual-aspect Attention mechanism significantly increases computational complexity, which may affect the performance and practicality of the proposed method. ’**Complexity of the Model**: The DACT arch, while innovative, introduces additional complexity compared to traditional models’, ’This complexity may lead to longer training times and increased computational resource requirements, which could be a barrier for practical applications in resource-constrained environments’ Yes Yes Yes
@@
**Insufficient Detailed Experimental Analysis**: The paper primarily provides quantitative results without a detailed experimental analysis, which limits the understanding of the findings. **Limited Discussion on Failure Cases**: While the paper discusses success and failure cases of existing methods, it may not provide sufficient insight into the specific conditions under which SCILL might fail, limiting the understanding of its robustness. Yes No Yes
@@
**Connection to Practical Methods**: There is an unknown connection between the theoretical findings and popular ensemble methods used in practice, which raises questions about the practical relevance of the work. **Unexplored Variants of Ensemble Methods**: The research does not explore other ensemble methods beyond those mentioned, such as boosting or stacking, which may have different theoretical properties and could provide additional insights into the interpolating regime No Yes Yes
@@
**Lack of Exploration of Alternative Methods**: The paper does not sufficiently explore or compare the performance of other methods like MFVI and SGLD in the context of covariate shift, which could provide a more comprehensive understanding of the problem. **Absence of Dataset Details**: The dataset section is marked as "nan," indicating a lack of information about the datasets used for experimentation’, ’This omission makes it difficult to assess the validity and applicability of the findings, as the choice of dataset can significantly influence the results No No No
Table 23: LLM as a Judge (GPT 4, Llama 3) and User to decide whether Ground Truth Limitation and LLM Generated Limitation have the same topic or not.
Note: Here GPT 4, Llama 3, and User decision regarding both limitation match or not.
@@
Ground Truth Limitations: Our goal is to provide comprehensive study of learning algorithms, including their hyperparameters. However, our study is limited on the variety of other design choices to restrict the experimental complexity. While we studied both synthetic and realistic image datasets, both these datasets are relatively simple with the same small number of generative factors and each of the factor follows uniform distribution. For learning algorithms, we focus on studying the inductive bias on the representation format while fixing the model arch design which can impact the results. Moreover, we did not study hyperparameters beyond those related to the latent representations. Specifically, we did not study how the type and configurations of the optimizer and the batch size would change the results; instead, we followed common setups in previous studies. '''   LLM Refined Limitations: ’1. Here are the limitations extracted from the provided limitation:’, ’2. . **Experimental Complexity**: The study is limited in the variety of design choices to reduce experimental complexity.’, ’3. . **Dataset Complexity**: The study utilizes relatively simple synthetic and realistic image datasets, which have a small number of generative factors.’, ’4. . **Generative Factors**: Each generative factor in the datasets follows a uniform distribution, limiting the diversity of the data.’, ’5. . **Model arch**: The focus is on studying the inductive bias related to the representation format while keeping the model arch design fixed, which may influence the results.’, ’ 6. . **Hyperparameter Scope**: The study does not explore hyperparameters beyond those related to latent representations.’, ’7. . **Optimizer and Batch Size**: The impact of different types and configurations of optimizers and batch sizes on the results is not examined, as the study adheres to common setups from previous research.’
@@
Ground Truth Limitations: While our work provides useful starting point for understanding student feedback,there are limitations to our work. Addressing these limitations will be an important area for future research. Comments may not reflect real student feedback. The comments in our dataset are from users who have chosen to post publicly on youtube. Addi- tionally,the comments may include features specific to this online education setting. Thus,the comments may reflect real student comments from these courses. There is selection bias in lecture sources. Sight includes lectures that may be drawn from the most successful offerings of that course. The instructional quality may not be representative of typical instruction. Thus,inferences drawn about the instruction should be interpreted with caution,as they might not generalize to other lecture settings. We analyze only english comments. We analyze only english comments because the lecture content is given in english and the authors are most comfortable with english. As result,our rubric may not capture the types of feedback from nonenglish students watching lectures taught in english. we annotate small subsample of the data to assess the validity of the automatic labels,we conduct diagnostic study on small,randomly selected subset of the dataset,comprising approximately of the comments. Our work aims to establish preliminary evaluation of the humanmodel agreement and model annotations,and further validation of the automatic labels is necessary. '''   LLM Refined Limitations: Here are the limitations extracted from the provided limitation: 1. **Comments May Not Reflect Real Student Feedback**: The dataset consists of comments from users who have chosen to post publicly on YouTube, which may not accurately represent genuine student feedback. 2. **Features Specific to Online Education Setting**: The comments may include characteristics that are unique to the online education context, potentially skewing the interpretation of student feedback. 3. **Selection Bias in Lecture Sources**: The lectures included in the analysis may be drawn from the most successful offerings of a course, leading to a selection bias that does not represent typical instructional quality. 4. **Generalizability of Instructional Quality**: Inferences about instructional quality based on the selected lectures should be made with caution, as they may not generalize to other lecture settings. 5. **Analysis Limited to English Comments**: The analysis is restricted to comments in English, which may overlook feedback from non-English speaking students who are watching lectures taught in English. 6. **Small Subsample for Validity Assessment**: The study annotates a small subsample of the data to assess the validity of automatic labels, which may not provide a comprehensive evaluation of the entire dataset. 7. **Need for Further Validation of Automatic Labels**: The work establishes a preliminary evaluation of human-model agreement and model annotations, indicating that further validation of the automatic labels is necessary.

### v2

@@ 8 Conclusion
We present a new approach for automatically extracting, generating, and evaluating limitations in scientific articles. Our method explores incorporating cited works, accommodating top sections of the entire paper, and integrating review feedback to capture perspectives beyond those of the original authors. To evaluate the effectiveness of our system, we introduce a granular text evaluation framework that breaks down limitations into more minor points and employs LLMs as a Judge for assessing alignment. Human review validates our extraction and LLM-as-Judge pipeline, showing strong agreement with expert judgments.
@@
In this work, we focused on venues in natural language processing (ACL papers from 2023-2024) and machine learning (NeurIPS papers 2021-2022), and Biology domain papers from PeerJ, which ensures high relevance and quality but insufficient for broader generalizability. While this scope allows us to benchmark the performance of LLMs in extracting limitations from well-structured scientific texts, we acknowledge that the findings may not generalize to papers from other fields, such as social sciences, physics, chemistry, or mathematics where writing conventions and limitation styles may differ.
Due to high API costs, we did not experiment with GPT-4 or GPT-4o; instead, we opted for GPT-4o Mini as a cost-effective alternative. While we incorporated OpenReview comments for NeurIPS papers, we could not find them for ACL papers. Furthermore, we relied on GPT-4o Mini as the evaluation judge. To evaluate the effectiveness of LLMs as both text extractors and judges, we conducted a human annotation study with 200 samples and only three annotators.
A key threat to validity is contamination bias, when evaluation examples (or close paraphrases) appear in a model’s training data, artificially inflating performance. To guard against this, we tested whether GPT-4o mini had been trained on our NeurIPS 2021–2022 dataset by providing only each paper’s title and prompting it to summarize the content and identify limitations. In every case, the model replied with a disclaimer indicating unfamiliarity with the specific work (e.g., “I am not familiar with the specific paper titled …”). This consistent outcome suggests the model lacked prior exposure to the full texts, supporting the integrity of our evaluation.
While we selected GPT-4o mini for text extraction, generation, and evaluation due to its superior performance, relying on a single LLM for these roles introduces several potential biases. We took specific steps to mitigate these risks: To counter self-validation bias, where the model might favor its own output, we cross-referenced its judgments with human evaluations and incorporated RAG. For positional bias, where the model may favor the first input when comparing texts, we swapped the input order to ensure consistent results. To reduce confirmation bias, the tendency to generate generic limitations, we used RAG to introduce more diverse evidence. Finally, to check for hallucinations, three human annotators verified that all extracted limitations were grounded in the source text. Although these strategies are crucial for improving reliability, we acknowledge that they do not completely eliminate these inherent biases.
For future work, we will expand our dataset to more diverse domains (e.g., bioinformatics, cognitive science) to test the cross-domain robustness of our models. We also plan to enhance our generation framework by exploring more advanced multi-agent and open-source LLMs via RAG. Finally, we will scale our human validation efforts with a larger, more diverse pool of expert annotators to enable a deeper and more reliable analysis.
@@
This research adheres to ACL ethical standards. All data, including research papers and OpenReview feedback, were sourced from public repositories in compliance with their usage policies and were not filtered based on discriminatory attributes. Our user study involved three computer science graduate students who participated voluntarily with no conflicts of interest.
We acknowledge and address inherent LLM risks, including biases from training corpora, confirmation bias toward “safe” limitations, fluency and verbosity biases favoring longer or well-written outputs, and self-validation bias when using the same model for multiple tasks. To mitigate these, we (1) ground all generations in source content and peer reviews via a RAG framework to improve factuality and reduce verbosity; (2) diversify our ground truth by incorporating human-authored OpenReview critiques; (3) use multiple models to break self-validation circularity; and (4) conduct parallel human evaluations to detect overconfidence and other model-specific biases. We recognize that further work is needed to rigorously quantify these issues and plan to investigate cross-domain robustness in future studies.
@@ A.2 Performance Measurement
We used three distinct prompts to check the topic-level similarity between ground truth limitations and LLM-generated limitations (Figure 5, Appendix). To overcome the position bias, we choose the consistent one.
@@
Prompt = ''' Here is the text containing extracted limitations. Please identify and list each limitation, ensuring that each one addresses a distinct topic or point. '''
Figure 3: Prompt to extract limitations from ground truth text.
Prompt = ''' You are a helpful, respectful, and honest assistant for generating limitations or shortcomings of a research paper. I am providing ’Abstract’, ’Introduction’, ’Related Work’, ’Methodology’, ’Experiment and Results’, ’Conclusion’, and other sections of a scientific paper alongside the related cited papers texts. Generate limitations based on these texts. '''
Figure 4: Prompt to generate limitations from Input and cited papers text.
@@
Figure 5: LLM as a Judge for each limitation. We use three distinct prompts to verify consistency.
@@
Model Complexity Concerns: - There is a question regarding whether the performance gains are due to increased model complexity rather than the proposed recursive mixing approach. Potential for Increased Complexity: Although the method is described as simple, the recursive nature of the approach may introduce complexity in implementation and understanding, particularly for practitioners who may not be familiar with the underlying concepts Yes Yes Yes
@@
Computational Complexity: The Dual-aspect Attention mechanism significantly increases computational complexity, which may affect the performance and practicality of the proposed method. ’**Complexity of the Model**: The DACT arch, while innovative, introduces additional complexity compared to traditional models’, ’This complexity may lead to longer training times and increased computational resource requirements, which could be a barrier for practical applications in resource-constrained environments’ Yes Yes Yes
@@
Insufficient Detailed Experimental Analysis: The paper primarily provides quantitative results without a detailed experimental analysis, which limits the understanding of the findings. Limited Discussion on Failure Cases: While the paper discusses success and failure cases of existing methods, it may not provide sufficient insight into the specific conditions under which SCILL might fail, limiting the understanding of its robustness. Yes No Yes
@@
Connection to Practical Methods: There is an unknown connection between the theoretical findings and popular ensemble methods used in practice, which raises questions about the practical relevance of the work. Unexplored Variants of Ensemble Methods: The research does not explore other ensemble methods beyond those mentioned, such as boosting or stacking, which may have different theoretical properties and could provide additional insights into the interpolating regime No Yes Yes
@@
Lack of Exploration of Alternative Methods: The paper does not sufficiently explore or compare the performance of other methods like MFVI and SGLD in the context of covariate shift, which could provide a more comprehensive understanding of the problem. Absence of Dataset Details: The dataset section is marked as "nan," indicating a lack of information about the datasets used for experimentation’, ’This omission makes it difficult to assess the validity and applicability of the findings, as the choice of dataset can significantly influence the results No No No
Table 9: Examples of Annotator, GPT 4o-mini, and LLama judgement on whether a generated limitation should be matched with a ground-truth limitation or not.
@@
Ground Truth Limitations: Our goal is to provide comprehensive study of learning algorithms, including their hyperparameters. However, our study is limited on the variety of other design choices to restrict the experimental complexity. While we studied both synthetic and realistic image datasets, both these datasets are relatively simple with the same small number of generative factors and each of the factor follows uniform distribution. For learning algorithms, we focus on studying the inductive bias on the representation format while fixing the model arch design which can impact the results. Moreover, we did not study hyperparameters beyond those related to the latent representations. Specifically, we did not study how the type and configurations of the optimizer and the batch size would change the results; instead, we followed common setups in previous studies. '''  LLM Extracted Limitations: 1. Here are the limitations extracted from the provided limitation:, 2. **Experimental Complexity**: The study is limited in the variety of design choices to reduce experimental complexity., 3. **Dataset Complexity**: The study utilizes relatively simple synthetic and realistic image datasets, which have a small number of generative factors., 4. **Generative Factors**: Each generative factor in the datasets follows a uniform distribution, limiting the diversity of the data., 5. **Model arch**: The focus is on studying the inductive bias related to the representation format while keeping the model arch design fixed, which may influence the results., 6. **Hyperparameter Scope**: The study does not explore hyperparameters beyond those related to latent representations., 7. **Optimizer and Batch Size**: The impact of different types and configurations of optimizers and batch sizes on the results is not examined, as the study adheres to common setups from previous research.
@@
Ground Truth Limitations: While our work provides useful starting point for understanding student feedback,there are limitations to our work. Addressing these limitations will be an important area for future research. Comments may not reflect real student feedback. The comments in our dataset are from users who have chosen to post publicly on youtube. Addi- tionally,the comments may include features specific to this online education setting. Thus,the comments may reflect real student comments from these courses. There is selection bias in lecture sources. Sight includes lectures that may be drawn from the most successful offerings of that course. The instructional quality may not be representative of typical instruction. Thus,inferences drawn about the instruction should be interpreted with caution,as they might not generalize to other lecture settings. We analyze only english comments. We analyze only english comments because the lecture content is given in english and the authors are most comfortable with english. As result,our rubric may not capture the types of feedback from nonenglish students watching lectures taught in english. we annotate small subsample of the data to assess the validity of the automatic labels,we conduct diagnostic study on small,randomly selected subset of the dataset,comprising approximately of the comments. Our work aims to establish preliminary evaluation of the humanmodel agreement and model annotations,and further validation of the automatic labels is necessary. '''  LLM Extracted Limitations: Here are the limitations extracted from the provided limitation: 1. **Comments May Not Reflect Real Student Feedback**: The dataset consists of comments from users who have chosen to post publicly on YouTube, which may not accurately represent genuine student feedback. 2. **Features Specific to Online Education Setting**: The comments may include characteristics that are unique to the online education context, potentially skewing the interpretation of student feedback. 3. **Selection Bias in Lecture Sources**: The lectures included in the analysis may be drawn from the most successful offerings of a course, leading to a selection bias that does not represent typical instructional quality. 4. **Generalizability of Instructional Quality**: Inferences about instructional quality based on the selected lectures should be made with caution, as they may not generalize to other lecture settings. 5. **Analysis Limited to English Comments**: The analysis is restricted to comments in English, which may overlook feedback from non-English speaking students who are watching lectures taught in English. 6. **Small Subsample for Validity Assessment**: The study annotates a small subsample of the data to assess the validity of automatic labels, which may not provide a comprehensive evaluation of the entire dataset. 7. **Need for Further Validation of Automatic Labels**: The work establishes a preliminary evaluation of human-model agreement and model annotations, indicating that further validation of the automatic labels is necessary.